# A Low-Rank ADMM Splitting Approach for Semidefinite Programming

### 14 Mar 2024 v1

@@ 5.4 The Matrix Completion Problems
The Matrix Completion Problem focuses on estimating the unobserved entries of a partially observed matrix and has a broad range of applications, including collaborative filtering, image inpainting, and system identification. Semidefinite programming is a well-known method for addressing matrix completion problems. However, its application in real-world practices, especially the large-scale cases, has been limited due to computational inefficiencies, as mentioned in Li et al. (2019). Despite this, recent advancements on SDP algorithms, such as those documented by Wang and Hu (2023), have demonstrated significant improvements over traditional SDP solvers, revealing the SDP approach’s potential. Additionally, recent contributions by Yalçin et al. (2023) indicate that the Burer-Monteiro factorization method is particularly advantageous in scenarios with limited observations and a small number of measurements compared with the traditional SDP approach. With interest, we test our method on matrix completion problems and observe that it yields promising outcomes.
@@
Given matrix $M\in\mathbb{R}^{n\times m}$ with observed entries in $\Omega\subseteq\{1,\ldots,n\}\times\{1,\ldots,m\}$, the problem is to find $Y$ such that $Y_{ij}=M_{ij}$ for any $(i,j)\in\Omega$, while rank($Y$) being minimized. The convex optimization problem can be formulated as follows:
@@
where $\|Y\|_{*}$ is the nuclear norm of $Y$. Further, the unclear norm minimization problem above can be formulated into a semidefinite programming problem as follows:
@@
$\min_{X\in\mathbb{S}^{(m+n)\times(m+n)}}\ \ \big{\langle}I,X\big{\rangle},\ \ \operatornamewithlimits{s.t.}\ \ \big{\langle}{\left[\begin{array}[]{cc}0_{m\times m}&E_{ij}^{\top}\\ E_{ij}&0_{n\times n}\end{array}\right]},X\big{\rangle}=2M_{ij},\ \ \forall(i,j)\in\Omega,X\succeq 0,$
@@
We benchmark our algorithm on a randomly generated set of matrix completion problems. The generation process follows the methodology outlined by Wang and Hu (2023).
As indicated by the results in Table 5, our solver consistently outperforms other solvers across a range of instance sizes. For instances larger than MC_10000, further evaluations of other solvers were not conducted due to the significantly long solving times. Notably, our solver is distinguished by its ability to successfully handle an exceptionally large-scale instance, featuring up to $15,694,167$ constraints and a $40000\times 40000$ positive semidefinite variable, in less than 6 minutes.
@@ 5.5 General SDP Benchmark Problems
In this section, we evaluate our method across a wider range of SDP problems from SDPLIB [^7] For the problem files and a detailed description, see http://euler.nmt.edu/~brian/sdplib/ and Hans Mittelmann’s SDP benchmarks [^8] For the problem files and a detailed description, see https://plato.asu.edu/ftp/sparse_sdp.html. SDPLIB offers relatively small-scale SDP instances, whereas Hans Mittelmann’s SDP benchmarks provides larger and more challenging SDP instances.
The comparative performance of the four solvers is detailed in Table 6. Our method demonstrates significant advancements over the Burer-Monteiro low-rank factorization approach (SDPLR) in the majority of tested instances. When compared with SDPNAL+, our solver falls short in only a few cases. Against COPT, our solver shows competitive performance, outperforming COPT in many instances while being outperformed in others. We need to remark that the results of LoRADS are based on a very elementary parameter tuning detailed in Section 5.1. For example, the problem "hand" can be solved in $5.81$ seconds if a more suitable $\rho_{max}$ and heuristic factor $h$ is applied, which reduces the solving time by a factor of 20.
@@
To further benchmark the performance of LoRADS, we employ the scaled shifted geometric means (SSGM) to analyze solving times for problems sourced from both SDPLIB and Hans Mittelmann’s benchmarks respectively. The SSGMs calculated for SDPLIB illurstrate LoRADS’s proficiency in solving small-scale SDP problems, whereas the SSGMs for Hans Mittelmann’s benchmarks assess the efficacy of LoRADS in addressing larger SDP challenges. For an observation $y\in\mathbb{R}^{n}_{+}$, the shifted geometric mean is calculated as follows:
@@ 6 Conclusion
It is important to remark that the development and tuning of LoRADS are still in their preliminary level. Future enhancements, such as the adaptive step sizes for the ADMM algorithm, more sophisticated warm-start strategies, intelligent automation for switching phases, and additional heuristic methods, are anticipated to significantly elevate the performance of LoRADS. Moreover, the current design of LoRADS is entirely a first-order framework, which makes it possible for the GPU-based implementation. With the integration of GPU’s parallel computing capabilities, it is promising that LoRADS may unlock its full potential for addressing super large-scale SDP challenges.
@@ Acknowledgments
The authors thank Wenzhi Gao from Standford University and Mingyu Lei, Jiayu Zhang from Shanghai Jiao Tong University for their insightful discussions and assistance. All authors are considered to have contributed equally to this work.
@@ A.1 Proof of Theorem 1
For notational simplicity, we will drop the superscript $\sigma$ in the constants $L_{f}^{\sigma},L^{\sigma}$ and $L_{0}^{\sigma}$ in the proof. Firstly, using the $\epsilon$-KKT condition (8) of (4), we have
@@
where the inequalities holds because $(\overline{U},\overline{V},\overline{\lambda})$ lies in the bounded set $\mathcal{B}_{\sigma}\big{(}0^{n\times r},0^{n\times r},0^{m}\big{)}$ and $\|\nabla f\|_{F}$ is bounded above by $L_{f}$ on $B_{\sigma^{2}}\big{(}0^{n\times n}\big{)}$. Next, adding up the first two equations in (8) and using the triangle inequality yield
$\|\nabla f(\overline{U}\overline{V}^{\top})\overline{V}+\nabla f(\overline{V}\overline{U}^{\top})\overline{U}+2\sum_{i=1}^{m}\overline{\lambda}_{i}A_{i}\widehat{U}\|_{F}\leq 2\epsilon$ (25)
@@
By using [Ortega and Rheinboldt, 2000, Proposition 3.2.12] and the Lipschitz continuity of $\nabla^{2}f$, we know
$\|\nabla f(X)-\nabla f(Y)-\nabla^{2}f(Y)(X-Y)\|\leq{L_{0}\over 2}\|X-Y\|_{F}^{2}$
@@
$\Big{\|}2\nabla f(\widehat{U}\widehat{U}^{\top})\widehat{U}+2\sum_{i=1}^{m}\overline{\lambda}_{i}A_{i}\widehat{U}\Big{\|}\leq 2\epsilon+\frac{(2L_{0}\sigma^{3}+3L\sigma)(\epsilon+L_{f}\sigma+\sqrt{s_{A}}\sigma^{2})^{2}}{2\gamma^{2}}.$
@@
$\leq\epsilon+\frac{\|\mathcal{A}\|_{2}(L_{f}\sigma+\sqrt{s_{A}}\sigma^{2})^{2}}{4\gamma^{2}}.$
@@ A.2 Proof of Proposition 1
$\begin{split}(\textrm{I})=&\mathcal{L}_{\rho}(U^{k+1},V^{k+1},{\lambda}^{k+1})-\mathcal{L}_{\rho}(U^{k+1},V^{k+1},{\lambda}^{k})\\ =&({\lambda}^{k+1}-{\lambda}^{k})^{\top}\big{(}{\cal A}(U^{k+1}(V^{k+1})^{\top})-b\big{)}\\ =&\rho\|{\cal A}(U^{k+1}(V^{k+1})^{\top})-b\|^{2}\\ =&\frac{1}{\rho}\|{\lambda}^{k+1}-{\lambda}^{k}\|^{2}.\end{split}$
@@ A.3 Proof of Proposition 2
For any given $\beta>0$, $S_{\beta}$ is closed because $f$ is a smooth convex function. Suppose $S_{\beta}$ is not empty.
(i) Let $\mu$ be the strongly convex modulus of $f(X)+\frac{\rho_{0}}{2}\|\mathcal{A}X-b\|^{2}$ and $X^{*}$ be its minimum point. It follows from the strongly convexity that
$f(UV^{T})+\frac{\rho_{0}}{2}\|\mathcal{A}(UV^{T})-b\|^{2}\geq f(X^{*})+\frac{\rho_{0}}{2}\|\mathcal{A}(X^{*})-b\|^{2}+\frac{\mu}{2}\|UV^{T}-X^{*}\|_{F}^{2},$
which implies $\{UV^{T}\,|\,(U,V)\in S_{\beta}\}$ is bounded. On the other hand, $\|U-V\|_{F}$ is bounded for any $(U,V)\in S_{\beta}$. Note that
$\|U\|_{F}^{2}+\|V\|_{F}^{2}=\|U-V\|_{F}^{2}+2{\rm trace}\,(UV^{T})$
@@
where the second inequality holds because the fact $\big{\langle}B_{1},B_{2}\big{\rangle}\geq-\|B_{1}\|_{F}\|B_{2}\|_{F}$ for any $B_{1},B_{2}$, the third inequality holds because of $\|B_{1}B_{2}\|_{F}\leq\|B_{1}\|_{2}\|B_{2}\|_{F}$, and the last inequlity holds because $\lambda_{\max}\|U\|_{F}\|V-U\|_{F}\leq\frac{\lambda_{\min}}{2}\|U\|_{F}^{2}+\frac{\lambda^{2}_{\max}}{2\lambda_{\min}}\|V-U\|_{F}^{2}$ and $\kappa_{c}=\frac{\lambda_{\max}}{\lambda_{\min}}$. For the second term in (30), we have
$\big{\langle}{y}^{*},\mathcal{A}(UV^{\top})\big{\rangle}=\big{\langle}{y}^{*},\mathcal{A}(UV^{\top})-b\big{\rangle}+\big{\langle}{y}^{*},b\big{\rangle}\geq-\frac{1}{2}\|{y}^{*}\|^{2}-\frac{1}{2}\|{\cal A}(UV^{\top})-b\|^{2}+\big{\langle}{y}^{*},b\big{\rangle}.$
@@
$f(UV^{\top})+\frac{\rho_{0}}{2}\|U-V\|_{F}^{2}+\frac{\rho_{0}}{2}\|{\cal A}(UV^{\top})-b\|^{2}$
$\geq$ $\frac{\lambda_{\min}}{2}\|U\|_{F}^{2}+\left(\frac{\rho_{0}}{2}-\frac{\lambda_{\max}\kappa_{c}}{2}\right)\|V-U\|_{F}^{2}+\left(\frac{\rho_{0}}{2}-\frac{1}{2}\right)\|{\cal A}(UV^{\top})-b\|^{2}-\frac{1}{2}\|{y}^{*}\|^{2}+\big{\langle}{y}^{*},b\big{\rangle}$
$\geq$ $\frac{\lambda_{\min}}{2}\|U\|_{F}^{2}-\frac{1}{2}\|{y}^{*}\|^{2}+\big{\langle}{y}^{*},b\big{\rangle}.$
@@
$\|U\|_{F}^{2}\leq\frac{1}{\lambda_{\min}}\left(\|{y}^{*}\|^{2}-2\big{\langle}{y}^{*},b\big{\rangle}+2\beta\right).$ (31)
@@ A.4 Proof of Lemma 1
We establish part (i) of this lemma by mathematical induction while proving part (ii)-(iv) automatically in the process. As $(U^{0},V^{0},\lambda^{0})$ is selected according to (19), part (i) naturally holds for $k=0$. Now, assuming part (i) holds for $k$-th iterate, we proceed to establish its validity for $k+1$. By this assumption and Proposition 1, we obtain
$\begin{split}\hat{\beta}&>L_{\rho}(U^{k},V^{k},{\lambda}^{k})+\frac{1}{\rho}\|\lambda^{k}\|^{2}\\ &\geq L_{\rho}(U^{k+1},V^{k+1},{\lambda}^{k})+\frac{1}{\rho}\|\lambda^{k}\|^{2}\\ &=f(U^{k+1}(V^{k+1})^{T})+\frac{\gamma}{2}\|U^{k+1}-V^{k+1}\|^{2}+\big{\langle}{\lambda}^{k},\mathcal{A}(U^{k+1}V^{k+1\top})-b\big{\rangle}+\frac{\rho}{2}\|\mathcal{A}(U^{k+1}V^{k+1\top})-b\|^{2}+{1\over\rho}\|\lambda^{k}\|^{2}\\ &\geq f(U^{k+1}(V^{k+1})^{T})+\frac{\gamma}{2}\|U^{k+1}-V^{k+1}\|^{2}-\frac{\|{\lambda}^{k}\|^{2}}{2(\rho-\rho_{0})}-\frac{(\rho-\rho_{0})}{2}\|\mathcal{A}(U^{k+1}V^{k+1\top})-b\|^{2}+{1\over\rho}\|\lambda^{k}\|^{2}\\ &\quad+\frac{\rho}{2}\|\mathcal{A}(U^{k+1}V^{k+1\top})-b\|^{2}\\ &\geq f(U^{k+1}(V^{k+1})^{T})+\frac{\rho_{0}}{2}\|U^{k+1}-V^{k+1}\|^{2}+\frac{\rho_{0}}{2}\|\mathcal{A}(U^{k+1}V^{k+1\top})-b\|^{2}+\left({1\over\rho}-{1\over 2(\rho-\rho_{0})}\right)\|{\lambda}^{k}\|^{2}\\ &\geq f(U^{k+1}(V^{k+1})^{T})+\frac{\rho_{0}}{2}\|U^{k+1}-V^{k+1}\|^{2}+\frac{\rho_{0}}{2}\|\mathcal{A}(U^{k+1}V^{k+1\top})-b\|^{2},\end{split}$
where the last inequality follows from $\rho>2\rho_{0}$. This implies $(U^{k+1},V^{k+1})\in S_{\hat{\beta}}$ and it thus holds from Assumption 2 that $\|U^{k+1}\|_{F}\leq\delta$ and $\|V_{F}\|^{k+1}\leq\delta$. Analogous to the proof above, we can also establish
$f(U^{k}(V^{k})^{T})+\frac{\rho_{0}}{2}\|U^{k}-V^{k}\|^{2}+\frac{\rho_{0}}{2}\|\mathcal{A}(U^{k}V^{k\top})-b\|^{2}\leq\hat{\beta}.$
@@
Note that $(U^{k+1},V^{k+1})\in S_{\hat{\beta}}$ and it follows from Assumption 3 that there exists some orthogonal matrix such that
$\sigma_{\min}\left(C(U^{k+1}R)\right)\geq\eta.$ (33)
Let $B_{k+1}=\nabla f(V^{k+1}(U^{k+1})^{\top})U^{k+1}R-\gamma(U^{k+1}R-V^{k+1}R)$, and $b_{k+1}=\operatorname{vec}(B_{k+1})\in\mathbb{R}^{nr}$. Let $C^{k+1}=C(U^{k+1}R)=[\operatorname{vec}{(}A_{1}U^{k+1}R),\cdots,\operatorname{vec}{(}A_{m}U^{k+1}R)]\in\mathbb{R}^{nr\times m}$. Then we can rewrite equation (32) as
$C^{k+1}\bm{\lambda}^{k+1}=-b^{k+1}.$ (34)
@@
$\begin{split}\eta\|\bm{\lambda}^{k+1}\|\leq&\|b^{k+1}\|=\|B^{k+1}\|_{F}\\ \leq&\|\nabla f(V^{k+1}(U^{k+1})^{\top})U^{k+1}\|+\gamma\|U^{k+1}-V^{k+1}\|\\ \leq&L_{f}\delta+2\gamma\delta.\end{split}$ (35)
@@
Then we can bound $\|\Delta{\lambda}^{k+1}\|$ by the following inequality
$\begin{split}\eta\|\Delta{\lambda}^{k+1}\|\leq\|C^{k}\Delta{\lambda}^{k+1}\|\leq\|b^{k+1}-b^{k}\|+\|C^{k+1}-C^{k}\|\|{\lambda}^{k+1}\|.\end{split}$ (36)
In what follows, we bound $\|b^{k+1}-b^{k}\|$ and $\|C^{k+1}-C^{k}\|$ separately
$\begin{split}\|b^{k+1}-b^{k}\|=&\|B^{k+1}-B^{k}\|_{F}\\ \leq&\|\nabla f\big{(}V^{k+1}(U^{k+1})^{T}\big{)}U^{k+1}R-\nabla f\big{(}V^{k}(U^{k})^{T}\big{)}U^{k}\|R+\gamma\|\Delta U^{k+1}R\|+\gamma\|\Delta V^{k+1}R\|\\ \leq&\big{(}L\delta^{2}+L_{f}+\gamma\big{)}\|\Delta U^{k+1}\|+\big{(}L\delta^{2}+\gamma\big{)}\|\Delta V^{k+1}\|,\end{split}$ (37)
@@
$\begin{split}\|C^{k+1}-C^{k}\|^{2}=\sum_{i=1}^{m}\|A_{i}\big{(}U^{k+1}-U^{k}\big{)}R\|^{2}\leq\sum_{i=1}^{m}\|A_{i}\|^{2}\|U^{k+1}-U^{k}\|^{2}=s_{A}\|\Delta U^{k+1}\|^{2},\end{split}$ (38)

### 25 Mar 2024 v2

@@ 5.4 The Matrix Completion Problem
The Matrix Completion Problem, which involves estimating the unobserved entries of a partially observed matrix, finds wide application in areas such as collaborative filtering, image inpainting, system identification, among others. SDP is a well-known method for addressing matrix completion problems. However, the use of conventional SDP solvers in matrix completion, particularly in large-scale instances, is hindered by computational inefficiencies (Li et al., 2019). Despite these limitations, recent progress in SDP algorithms, highlighted by Wang and Hu (2023), has demonstrated significant improvements over traditional SDP solvers. Additionally, recent studies by Yalçin et al. (2023) have shown that the Burer-Monteiro factorization method is particularly advantageous in scenarios with limited observations and a small number of measurements. Motivated by these efforts, we test our method on matrix completion problems and demonstrate its effectiveness.
@@
Given a matrix $M\in\mathbb{R}^{n\times m}$ with observed entries in $\Omega\subseteq\{1,\ldots,n\}\times\{1,\ldots,m\}$, the problem is to find $Y$ such that $Y_{ij}=M_{ij}$ for any $(i,j)\in\Omega$, while rank($Y$) being minimized. The convex optimization problem can be formulated as follows:
@@
where $\|Y\|_{*}$ is the nuclear norm of $Y$. Further, the nuclear norm minimization problem above can be formulated into a semidefinite programming problem as follows:
@@
$\min_{X\in\mathbb{S}^{(m+n)\times(m+n)}}\ \ \big{\langle}I,X\big{\rangle},\ \ \operatornamewithlimits{s.t.}\ \ \bigg{\langle}\left[\begin{array}[]{cc}0_{m\times m}&E_{ij}^{\top}\\ E_{ij}&0_{n\times n}\end{array}\right],{X}\bigg{\rangle}=2M_{ij},\ \ \forall(i,j)\in\Omega,X\succeq 0,$
@@
  • 1
    For instances beyond MC_10000, we only evaluate LoRADS and use "-" for other solvers as placeholders.
We benchmark our algorithm on a set of randomly generated matrix completion problems, adopting the generating process described by Wang and Hu (2023). The results, as depicted in Table 5, demonstrate that LoRADS consistently outperforms other solvers across a range of instance sizes. For instances larger than MC_10000, further evaluations of the compared solvers were not conducted due to time and computation resource limit. Notably, LoRADS successfully solves an exceptionally large-scale instance with $15,694,167$ constraints and a $40000\times 40000$ matrix variable in less than 6 minutes.
@@ 5.5 General SDP Benchmark Problems
In this section, we evaluate our method across a wider range of SDP problems from SDPLIB [^8] For the problem files and a detailed description, see http://euler.nmt.edu/~brian/sdplib/. and Hans Mittelmann’s SDP benchmarks [^9] For the problem files and a detailed description, see https://plato.asu.edu/ftp/sparse_sdp.html.. SDPLIB offers relatively small-scale SDP instances, while Hans Mittelmann’s SDP benchmarks provide larger and more challenging SDP instances.
The comparative performance of the four solvers is detailed in Table 6. Our method exhibits considerable improvements over SDPLR across most tested instances. In comparison with SDPNAL+, our solver falls short in only a few cases. Against COPT, our solver shows competitive performance, surpassing COPT in several instances, albeit falling short in others. We need to remark that the results of LoRADS are based on a very elementary parameter tuning detailed in Section 5.1. For example, the problem "hand" can be solved in $5.81$ seconds if a more suitable $\rho_{max}$ and heuristic factor $h$ is applied, potentially reducing the solving time by a factor of 20.
@@
To further benchmark the performance of LoRADS, we employ the scaled shifted geometric means (SSGM) to analyze solving times for problems sourced from both SDPLIB and Hans Mittelmann’s benchmarks respectively. The SSGMs calculated for SDPLIB illustrate LoRADS’s proficiency in solving small-scale SDP problems, whereas the SSGMs for Hans Mittelmann’s benchmarks assess the efficacy of LoRADS in addressing larger SDP challenges. For an observation $y\in\mathbb{R}^{n}_{+}$, the shifted geometric mean is calculated as follows:
@@ 6 Conclusion
It is important to remark that the development and tuning of LoRADS are still at their preliminary level. Future enhancements, such as the adaptive step sizes for the ADMM algorithm, more sophisticated warm-start strategies, intelligent automation for switching phases, and additional heuristic methods, are anticipated to significantly elevate the performance of LoRADS. Moreover, the current design of LoRADS is entirely a first-order framework, which makes it possible for the GPU-based implementation. With the integration of GPU’s parallel computing capabilities, it is promising that LoRADS may unlock its full potential for addressing super large-scale SDP challenges.
@@ Acknowledgments
The authors thank Wenzhi Gao from Standford University and Mingyu Lei and Jiayu Zhang from Shanghai Jiao Tong University for their insightful discussions and assistance.
@@ A.1 Proof of Theorem 1
Proof For notation simplicity, we will drop the superscript $\sigma$ in the constants $L_{f}^{\sigma},L^{\sigma}$ and $L_{0}^{\sigma}$ in the proof. Firstly, using the $\epsilon$-KKT condition (8) of (4), we have
@@
where the inequalities holds because $(\overline{U},\overline{V},\overline{\lambda})$ lies in the bounded set $\mathcal{B}_{\sigma}\big{(}0^{n\times r},0^{n\times r},0^{m}\big{)}$, $\|\nabla f\|_{F}$ is bounded above by $L_{f}$ on $B_{\sigma^{2}}\big{(}0^{n\times n}\big{)}$ and
$\left\|\sum_{i=1}^{m}\overline{\lambda}_{i}A_{i}\overline{U}\right\|_{F}\leq\sum_{i=1}^{m}|\overline{\lambda}_{i}|\cdot\|A_{i}\|_{2}\cdot\|\overline{U}\|_{F}\leq\sqrt{s_{A}}\|\overline{{\lambda}}\|_{2}\cdot\|\overline{U}\|_{F}$
where the first inequality holds because of triangle inequality and the second inequality because of the Cauchy–Schwarz inequality. Next, adding up the first two equations in (8) and using the triangle inequality yield
$\Big{\|}\nabla f(\overline{U}\overline{V}^{\top})\overline{V}+\nabla f(\overline{V}\overline{U}^{\top})\overline{U}+2\sum_{i=1}^{m}\overline{\lambda}_{i}A_{i}\widehat{U}\Big{\|}_{F}\leq 2\epsilon$ (25)
@@
Using [Ortega and Rheinboldt, 2000, Proposition 3.2.12] and the Lipschitz continuity of $\nabla^{2}f$, we know
$\|\nabla f(X)-\nabla f(Y)-\nabla^{2}f(Y)(X-Y)\|_{F}\leq{L_{0}\over 2}\|X-Y\|_{F}^{2},$
@@
$\Big{\|}2\nabla f(\widehat{U}\widehat{U}^{\top})\widehat{U}+2\sum_{i=1}^{m}\overline{\lambda}_{i}A_{i}\widehat{U}\Big{\|}_{F}\leq 2\epsilon+\frac{(2L_{0}\sigma^{3}+3L\sigma)(\epsilon+L_{f}\sigma+\sqrt{s_{A}}\sigma^{2})^{2}}{2\gamma^{2}}.$
@@
$\leq\epsilon+\frac{\|\mathcal{A}\|_{2}(\epsilon+L_{f}\sigma+\sqrt{s_{A}}\sigma^{2})^{2}}{4\gamma^{2}}.$
@@ A.2 Proof of Proposition 1
$\begin{split}(\textrm{I})=&\mathcal{L}_{\rho}(U^{k+1},V^{k+1},{\lambda}^{k+1})-\mathcal{L}_{\rho}(U^{k+1},V^{k+1},{\lambda}^{k})\\ =&({\lambda}^{k+1}-{\lambda}^{k})^{\top}\big{(}{\cal A}(U^{k+1}(V^{k+1})^{\top})-b\big{)}\\ =&\rho\|{\cal A}(U^{k+1}(V^{k+1})^{\top})-b\|_{2}^{2}\\ =&\frac{1}{\rho}\|{\lambda}^{k+1}-{\lambda}^{k}\|_{2}^{2}.\end{split}$
@@ A.3 Proof of Proposition 2
Proof For any given $\beta>0$, $S_{\beta}$ is closed because $f$ is a smooth convex function. Suppose $S_{\beta}$ is not empty.
(i) Let $\mu$ be the strongly convex modulus of $f(X)+\frac{\rho_{0}}{2}\|\mathcal{A}(X)-b\|_{2}^{2}$ and $X^{*}$ be its minimum point. It follows from the strongly convexity that
$f(UV^{\top})+\frac{\rho_{0}}{2}\|\mathcal{A}(UV^{\top})-b\|_{2}^{2}\geq f(X^{*})+\frac{\rho_{0}}{2}\|\mathcal{A}(X^{*})-b\|_{2}^{2}+\frac{\mu}{2}\|UV^{\top}-X^{*}\|_{F}^{2},$
which implies $\{UV^{\top}\,|\,(U,V)\in S_{\beta}\}$ is bounded. On the other hand, $\|U-V\|_{F}$ is bounded for any $(U,V)\in S_{\beta}$. Note that
$\|U\|_{F}^{2}+\|V\|_{F}^{2}=\|U-V\|_{F}^{2}+2\big{\langle}U,V\big{\rangle}$
@@
where the second inequality holds because the fact $\big{\langle}B_{1},B_{2}\big{\rangle}\geq-\|B_{1}\|_{F}\|B_{2}\|_{F}$ for any $B_{1},B_{2}$, the third inequality holds because of $\|B_{1}B_{2}\|_{F}\leq\|B_{1}\|_{F}\|B_{2}\|_{F}$, and the last inequlity holds because $\lambda_{\max}\|U\|_{F}\|V-U\|_{F}\leq\frac{\lambda_{\min}}{2}\|U\|_{F}^{2}+\frac{\lambda^{2}_{\max}}{2\lambda_{\min}}\|V-U\|_{F}^{2}$ and $\kappa_{c}=\frac{\lambda_{\max}}{\lambda_{\min}}$. For the second term in (30), we have
$\big{\langle}{y}^{*},\mathcal{A}(UV^{\top})\big{\rangle}=\big{\langle}{y}^{*},\mathcal{A}(UV^{\top})-b\big{\rangle}+\big{\langle}{y}^{*},b\big{\rangle}\geq-\frac{1}{2}\|{y}^{*}\|_{2}^{2}-\frac{1}{2}\|{\cal A}(UV^{\top})-b\|_{2}^{2}+\big{\langle}{y}^{*},b\big{\rangle}.$
@@
$f(UV^{\top})+\frac{\rho_{0}}{2}\|U-V\|_{F}^{2}+\frac{\rho_{0}}{2}\|{\cal A}(UV^{\top})-b\|_{2}^{2}$
$\geq$ $\frac{\lambda_{\min}}{2}\|U\|_{F}^{2}+\left(\frac{\rho_{0}}{2}-\frac{\lambda_{\max}\kappa_{c}}{2}\right)\|V-U\|_{F}^{2}+\left(\frac{\rho_{0}}{2}-\frac{1}{2}\right)\|{\cal A}(UV^{\top})-b\|_{2}^{2}-\frac{1}{2}\|{y}^{*}\|_{2}^{2}+\big{\langle}{y}^{*},b\big{\rangle}$
$\geq$ $\frac{\lambda_{\min}}{2}\|U\|_{F}^{2}-\frac{1}{2}\|{y}^{*}\|_{2}^{2}+\big{\langle}{y}^{*},b\big{\rangle}.$
@@
$\|U\|_{F}^{2}\leq\frac{1}{\lambda_{\min}}\left(\|{y}^{*}\|_{2}^{2}-2\big{\langle}{y}^{*},b\big{\rangle}+2\beta\right).$ (31)
@@ A.4 Proof of Lemma 1
Proof We establish part (i) of this lemma by mathematical induction while proving part (ii)-(iv) automatically in the process. As $(U^{0},V^{0},\lambda^{0})$ is selected according to (19), part (i) naturally holds for $k=0$. Now, assuming part (i) holds for $k$-th iterate, we proceed to establish its validity for $k+1$. By this assumption and Proposition 1, we obtain
$\begin{split}\hat{\beta}&>L_{\rho}(U^{k},V^{k},{\lambda}^{k})+\frac{1}{\rho}\|\lambda^{k}\|_{2}^{2}\\ &\geq L_{\rho}(U^{k+1},V^{k+1},{\lambda}^{k})+\frac{1}{\rho}\|\lambda^{k}\|_{2}^{2}\\ &=f(U^{k+1}(V^{k+1})^{T})+\frac{\gamma}{2}\|U^{k+1}-V^{k+1}\|_{F}^{2}+\big{\langle}{\lambda}^{k},\mathcal{A}(U^{k+1}V^{k+1\top})-b\big{\rangle}+\frac{\rho}{2}\|\mathcal{A}(U^{k+1}V^{k+1\top})-b\|_{2}^{2}+{1\over\rho}\|\lambda^{k}\|_{2}^{2}\\ &\geq f(U^{k+1}(V^{k+1})^{T})+\frac{\gamma}{2}\|U^{k+1}-V^{k+1}\|_{F}^{2}-\frac{\|{\lambda}^{k}\|_{2}^{2}}{2(\rho-\rho_{0})}-\frac{(\rho-\rho_{0})}{2}\|\mathcal{A}(U^{k+1}V^{k+1\top})-b\|_{2}^{2}+{1\over\rho}\|\lambda^{k}\|_{2}^{2}\\ &\quad+\frac{\rho}{2}\|\mathcal{A}(U^{k+1}V^{k+1\top})-b\|_{2}^{2}\\ &\geq f(U^{k+1}(V^{k+1})^{T})+\frac{\rho_{0}}{2}\|U^{k+1}-V^{k+1}\|_{F}^{2}+\frac{\rho_{0}}{2}\|\mathcal{A}(U^{k+1}V^{k+1\top})-b\|_{2}^{2}+\left({1\over\rho}-{1\over 2(\rho-\rho_{0})}\right)\|{\lambda}^{k}\|_{2}^{2}\\ &\geq f(U^{k+1}(V^{k+1})^{T})+\frac{\rho_{0}}{2}\|U^{k+1}-V^{k+1}\|_{F}^{2}+\frac{\rho_{0}}{2}\|\mathcal{A}(U^{k+1}V^{k+1\top})-b\|_{2}^{2},\end{split}$
where the last inequality follows from $\rho>2\rho_{0}$. This implies $(U^{k+1},V^{k+1})\in S_{\hat{\beta}}$ and it thus holds from Assumption 2 that $\|U^{k+1}\|_{F}\leq\delta$ and $\|V^{k+1}\|_{F}\leq\delta$. Analogous to the proof above, we can also establish
$f(U^{k}(V^{k})^{T})+\frac{\rho_{0}}{2}\|U^{k}-V^{k}\|_{F}^{2}+\frac{\rho_{0}}{2}\|\mathcal{A}(U^{k}V^{k\top})-b\|_{2}^{2}\leq\hat{\beta}.$
@@
Note that $(U^{k+1},V^{k+1})\in S_{\hat{\beta}}$ and it follows from Assumption 3 that
$\sigma_{\min}\left(\mathcal{C}(U^{k+1})\right)\geq\eta.$ (33)
Let $B_{k+1}=\nabla f(V^{k+1}(U^{k+1})^{\top})U^{k+1}-\gamma(U^{k+1}-V^{k+1})$, and $b_{k+1}=\operatorname{vec}(B_{k+1})\in\mathbb{R}^{nr}$. Let $C^{k+1}=\mathcal{C}(U^{k+1})=[\operatorname{vec}{(}A_{1}U^{k+1}),\cdots,\operatorname{vec}{(}A_{m}U^{k+1})]\in\mathbb{R}^{nr\times m}$. Then we can rewrite equation (32) as
$C^{k+1}{\lambda}^{k+1}=-b^{k+1}.$ (34)
@@
$\begin{split}\eta\|{\lambda}^{k+1}\|_{2}\leq&\|b^{k+1}\|_{2}=\|B^{k+1}\|_{F}\\ \leq&\|\nabla f(V^{k+1}(U^{k+1})^{\top})U^{k+1}\|_{F}+\gamma\|U^{k+1}-V^{k+1}\|_{F}\\ \leq&L_{f}\delta+2\gamma\delta.\end{split}$ (35)
@@
Then we can bound $\|\Delta{\lambda}^{k+1}\|_{2}$ by the following inequality
$\begin{split}\eta\|\Delta{\lambda}^{k+1}\|_{2}\leq\|C^{k}\Delta{\lambda}^{k+1}\|_{2}\leq\|b^{k+1}-b^{k}\|_{2}+\|C^{k+1}-C^{k}\|_{F}\|{\lambda}^{k+1}\|_{2}.\end{split}$ (36)
In what follows, we bound $\|b^{k+1}-b^{k}\|_{2}$ and $\|C^{k+1}-C^{k}\|_{F}$ separately
$\begin{split}\|b^{k+1}-b^{k}\|_{2}=&\|B^{k+1}-B^{k}\|_{F}\\ \leq&\|\nabla f\big{(}V^{k+1}(U^{k+1})^{T}\big{)}U^{k+1}-\nabla f\big{(}V^{k}(U^{k})^{T}\big{)}U^{k}\|_{F}+\gamma\|\Delta U^{k+1}\|_{F}+\gamma\|\Delta V^{k+1}\|_{F}\\ \leq&\big{(}L\delta^{2}+L_{f}+\gamma\big{)}\|\Delta U^{k+1}\|_{F}+\big{(}L\delta^{2}+\gamma\big{)}\|\Delta V^{k+1}\|_{F},\end{split}$ (37)
@@
$\begin{split}\|C^{k+1}-C^{k}\|_{F}^{2}=\sum_{i=1}^{m}\|A_{i}\big{(}U^{k+1}-U^{k}\big{)}\|_{F}^{2}\leq\sum_{i=1}^{m}\|A_{i}\|_{F}^{2}\|U^{k+1}-U^{k}\|_{F}^{2}=s_{A}\|\Delta U^{k+1}\|_{F}^{2},\end{split}$ (38)